OpenStudy (anonymous):

g(x)=cos^2(pi/8 x)+ sin (pi/4 x), then g(6)

OpenStudy (tylerd):

what are you trying to solve for exactly?

OpenStudy (tylerd):

g(6) ud just plug in 6 for x right?

OpenStudy (anonymous):

Yea just plug it in

OpenStudy (tylerd):

\[g(6)=\cos^2(\frac{ 6\pi }{ 8 })+\sin(\frac{ 6\pi }{ 4 })\]

OpenStudy (tylerd):

are you trying to simply this out any further?

OpenStudy (anonymous):

Yes

OpenStudy (tylerd):

ill have to look up the identities

OpenStudy (tylerd):

which identities have you recently gone over?

OpenStudy (anonymous):

No we haven't gone over any identities yet, this is just a problem from my summer calculus packet.

OpenStudy (tylerd):

you can do all kinds of things to it but

OpenStudy (tylerd):

unless your trying to find the reference angle im not sure what exactly we should do

OpenStudy (tylerd):

oh

OpenStudy (anonymous):

All I know you plug in the given but I don't know how the answer is suppose to look

OpenStudy (tylerd):

\[\cos^2(\frac{ 6\pi }{ 8 })+\sin(\frac{ 6\pi }{ 4 })= ?\]

OpenStudy (lυἶცἶ0210):

That's right, continue.

OpenStudy (tylerd):

well it looks like

OpenStudy (tylerd):

the theta in the cos^2 is half the size of the sin theta

OpenStudy (lυἶცἶ0210):

Try this: \(\Large \frac{1}{2}(1+cos~2(\frac{3\pi}{4}))+sin(\frac{3\pi}{2})\) Simplier now?

OpenStudy (lυἶცἶ0210):

It'll reduce to: \(\Large \frac{1}{2}(1+0)+(-1)\)

OpenStudy (lυἶცἶ0210):

Make sense?

OpenStudy (tylerd):

you used product to sum?

OpenStudy (anonymous):

Yea I think I get it

OpenStudy (lυἶცἶ0210):

I used the half angle formulas: \(\Large \frac{1}{2}(1+cos(2u))\) I find it easier to factor out the 1/2 right away

OpenStudy (lυἶცἶ0210):

^ That one being for \(cos^2x\)

OpenStudy (tylerd):

do you have all the identities memorized?

OpenStudy (lυἶცἶ0210):

Haha, not all of them, but with practice some just stick to you. Like the half angles are important later on.